OpenStudy (anonymous):

Quadratic Formula. y=-x^2-2

OpenStudy (anonymous):

x=(-b+/-radical(b^2-4ac)/2a

OpenStudy (anonymous):

yes but how do you plug in the numbers???

OpenStudy (anonymous):

\[x=-b \pm \sqrt{b^2-4ac}/2a\]

OpenStudy (anonymous):

No I mean how do I get the a, b, and c values????

OpenStudy (anonymous):

a=?,b=?,c=?????

OpenStudy (anonymous):

yeah.

OpenStudy (agent0smith):

Compare \[\Large y=-x^2-2\]to\[\Large y = ax^2 + bx + c\]and see if you can identify the values of a, b and c.

OpenStudy (anonymous):

Thanks.